-draw rectangle 27). Find the volume of the solid: y=x2-3 and y=15-x& around x= 6 - using Shell Method:
Answer #1

Given, I crista Chilstu) (6,15-12) ROM?+3) (no) nel hlu) = 15-n²-n²-3 = 12-242. R(u) = 6-u 56 V z 27 S (12-24) (6-4) du V6 =